Using the point-slpe formula, what is the equation of the line that passes throug the point (1,5) and (0,0)?
Answer by mananth(16949)

1,5) and (0,0)
the slope of line = m = y2-y1 / x2-x1
m= 0-5 / 0-1
=-5/-1
m=5 the slope.
..
to find the intercept plug the value of slope and any point above.
y= mx +b
0=5*0+b
so b = 0
..
The equation will ne y = 5x